Jen and Larry’s Frozen Yogurt Company

     In 2019, Jennifer (Jen) Liu and Larry Mestas founded Jean and Larry’s Frozen Yogurt Company, which was based on the idea of applying the microbrew or microbatch strategy to the production and sale of frozen yogurt. Jen and Larry began producing small quantities of unique flavors and blends in limited editions. Revenues were $600,000 in 2019 and were estimated to be $1.2 million in 2020.

     Because Jen and Larry were selling premium frozen yogurt containing premium ingredients, each small cup of yogurt sold for $3, and the cost of producing the frozen yogurt averaged $1.50 per cup. Administrative expenses, including Jen and Larry’s salary and expenses for an accountant and two other administrative staff, were estimated at $180,000 in 2020. Marketing expenses, largely in the form of behind-the-counter workers, in-store posters, and advertising in local newspapers, were projected to be $200,000 in 2020.

     An investment in bricks and mortar was necessary to make and sell the yogurt. Initial specialty equipment and the renovation of an old warehouse building in lower downtown (known as LoDo) occurred at the beginning of 2019. Additional equipment needed to make the amount of yogurt forecasted to be sold in 2020 was purchased at the beginning of 2020. As a result, depreciation expenses were expected to be $50,000 in 2020. Interest expenses were estimated at $15,000 in 2020. The average tax rate was expected to be 25% of taxable income.

  1. Calculate the dollar amount of NOPAT if Jen and Larry’s venture achieves the forecasted $1.2 million in sales in 2020. What would NOPAT be as a percent of sales?
  2. Calculate the NOPAT breakeven point for 2020 in terms of NOPAT breakeven revenues for Jen and Larry’s venture. How many cups of frozen yogurt would have to be sold to reach NOPAT breakeven?
Answer #1

Statement Showing Calculation of Net operating Income

Particulars Amount
Estimated Revenue 1,200,000
Less: Cost of Goods Sold(1,200,000/$3 per unit)*$1.5 per unit (600,000)
Gross Profit 600,000
Less:Salary (180,000)
Less:Advertisement Expense (200,000)
Less:Depreciation (50,000)
Net Operating Profit 170,000

a)

Calculation of NOPAT

NOPAT= Operating Profit* (1-Tax)

= 170,000*(1-0.25)

= 127,500

NOPAT be as percentage of Profit = NOPAT/Estimated Revenue

= 127,500/1,200,000

= 10.625%

b)

Calculation of NOPAT Break even point

NOPAT Break even point Revenue = Total Operating Cost / gross profit ratio

=(180,000+200,000+50,000)/0.5

=$860,000

Number of Units to be sold to achieve NOPAT BEP= BEP Revenue/ Unit Selling price

= 860,000/3

=286,667 units